# Dataforge: A Data Agent Platform for Autonomous Data Engineering

### v1

@@ III-B Quantitative Evaluations
TABLE I: Downstream performance comparison on regression and classification datasets. For regression, we report 1-RAE (Relative Absolute Error), and for classification, F1-score (%).
@@
FeatLLM [han2024large] 76.35% 93.62% 95.38% 89.66% 72.62% 81.17% 0.6477 0.4597 0.5587
CAAFE [hollmann2023large] 59.92% 91.41% 92.84% 79.86% 66.10% 81.74% N/A N/A N/A
AutoFeat [horn2019autofeat] 74.86% 93.29% 93.37% 80.86% 69.08% 82.54% 0.6329 0.4407 0.5746
OpenFE [zhang2023openfe] 74.50% 93.44% 93.37% 80.86% 69.09% 83.05% 0.6311 0.4402 0.5746
ELLM-FT [gong2025evolutionary] 76.39% 93.17% 96.01% 89.66% 74.80% 82.70% 0.6328 0.4734 0.6174
@@
RL-Policy-1 (GRFG [wang2022group]) 68.29% 68.29% 93.16% 75.39% 69.24% 79.81% 0.5768 0.4562 0.5587
RL-Policy-2 (MOAT [wang2023reinforcement]) 72.44% 72.44% 95.69% 80.73% 73.02% 81.17% 0.6251 0.4734 0.5967
DataAgent 79.60% 94.41% 97.14% 90.14% 76.98% 84.64% 0.7849 0.7219 0.7492
Study of Reliability and Automation. In practice, Dataforge showed stable behavior. It finished all nine datasets without interruption and usually converged within two action retries on average. Compared with Llama 3.1 8B and Llama 3 8B, which often fail to reach valid results at a 3–5% rate, the agent maintained smooth execution across runs. The average score reached 0.783, slightly higher than both baselines, and no invalid operations were observed. This suggests that the built-in routing and grounding steps make the workflow more predictable and robust for continuous operation.
Study of Efficiency Trade-offs. Compared with reinforcement-learning-based methods, Dataforge runs efficiently without any training cost. RL-Policy-1 and RL-Policy-2 spent over 1,400 s and 650 s on training, respectively, while the agent required none. During inference, the agent finished each workflow in about 3.9 s, with only two calls on average, whereas the RL-based baselines needed 13–22 calls per task. Even with this lightweight process, the agent still achieved the best performance (0.7832). These results show that the system keeps both accuracy and responsiveness, avoiding the heavy overhead of reinforcement-style optimization.
III-C Toward Accurate and Efficient Heart-Disease Detection
Figure 4: Dataforge Dealing with Heart-Disease Detection.
@@ III-C Toward Accurate and Efficient Heart-Disease Detection
Throughout this process (Fig. 4), no manual intervention or algorithmic knowledge is needed. The workflow completes in about 100 seconds and generates a concise report and a new dataset that achieves higher diagnostic accuracy with fewer features, reducing the total number of variables from 44 to 20 while improving the predictive score from 0.772 to 0.840. This demonstrates that DataForge can serve as an intelligent assistant for cardiology, automatically transforming complex measurements into a cleaner, smaller, and more reliable dataset for accurate and efficient heart-disease detection.

### v2

@@ 3.6. Pseudo-code of Dataforge
Table 1. Downstream performance comparison on regression and classification datasets. For regression, we report 1-RAE (Relative Absolute Error), and for classification, F1-score (%).
@@
FeatLLM (Han et al., 2024) 76.35% 93.62% 95.38% 89.66% 72.62% 81.17% 0.6477 0.4597 0.5587
CAAFE (Hollmann et al., 2023) 59.92% 91.41% 92.84% 79.86% 66.10% 81.74% N/A N/A N/A
AutoFeat (Horn et al., 2019) 74.86% 93.29% 93.37% 80.86% 69.08% 82.54% 0.6329 0.4407 0.5746
OpenFE (Zhang et al., 2023) 74.50% 93.44% 93.37% 80.86% 69.09% 83.05% 0.6311 0.4402 0.5746
ELLM-FT (Gong et al., 2025c) 76.39% 93.17% 96.01% 89.66% 74.80% 82.70% 0.6328 0.4734 0.6174
@@
RL-Policy-1 (GRFG (Wang et al., 2022)) 68.29% 68.29% 93.16% 75.39% 69.24% 79.81% 0.5768 0.4562 0.5587
RL-Policy-2 (MOAT (Wang et al., 2023)) 72.44% 72.44% 95.69% 80.73% 73.02% 81.17% 0.6251 0.4734 0.5967
Dataforge 79.60% 94.41% 97.14% 90.14% 76.98% 84.64% 0.7849 0.7219 0.7492
4. Empirical Study
4.1. Setup
Datasets and protocol. We evaluate on diverse tabular benchmarks from UCIrvine (Public, 2023c), CPLM (Public, 2023a), Kaggle (Howard, 2023), and OpenML (Public, 2023b), covering both classification and regression (statistics in Table 1). All methods are evaluated with the same 5-fold cross-validation protocol, and we report the average score across folds.
Downstream evaluator and metrics. We assess feature engineering utility by training a fixed downstream model on the dataset with identical hyperparameters for all methods. For classification, we report macro-F1; for regression, we report $1$-RAE:
$1\text{-RAE}=1-\frac{\|\bm{y}_{pred}-\bm{y}_{real}\|_{1}}{\|\bm{y}_{real}-\bar{\bm{y}}_{real}\|_{1}}.$
Operator set. All applicable methods share the same operator set $\mathcal{O}=\mathcal{O}_{1}\cup\mathcal{O}_{2}$, where $\mathcal{O}_{1}=\{\texttt{sqrt},\texttt{square},\texttt{cube},\texttt{reciprocal},\texttt{log},\\ \texttt{sin},\texttt{cos},\texttt{tanh},\texttt{sigmoid},\texttt{standard},\texttt{normalize},\texttt{quantile}\}$ and $\mathcal{O}_{2}=\{+,-,\times,\div\}$.
LLM backbones. We use Llama-3.2-3B as our base LLM for the agent. We use a unified generation protocol with $\texttt{temperature}=0.7$, $\texttt{top\_p}=0.9$, and $\texttt{max\_new\_tokens}=500$.
Baselines and implementation. We compare against: (i) no/linear baselines, including Original, RDG, PCA (Maćkiewicz and Ratajczak, 1993), and LDA; (ii) classical automated feature engineering and search methods, including ERG, AFAT (Horn et al., 2020), AutoFeat (Horn et al., 2019), NFS (Chen et al., 2019), TTG (Khurana et al., 2018), and OpenFE (Zhang et al., 2023); (iii) RL-based policies, including GRFG (Wang et al., 2022) and MOAT (Wang et al., 2023); and (iv) LLM-based approaches, including FeatLLM (Han et al., 2024), CAAFE (Hollmann et al., 2023), ELLM-FT (Gong et al., 2025c), as well as a Pure LLM variant.
Configurations. All experiments were conducted on the Ubuntu 22.04.3 LTS operating system, with a 13th-generation Intel(R) Core(TM) i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU. The experiments were conducted using Python 3.11.5 and PyTorch 2.0.1.
4.2. Study of Overall Performances.
Table 1 reports the downstream utility of different feature-engineering approaches on six classification and three regression benchmarks. We compare Dataforge with representative baselines from three families: classical automated feature engineering/search, RL-based policies, and LLM-based methods, as well as the Original features. The goal of this experiment is to evaluate whether an agentic, fully automated pipeline can consistently improve tabular representations across diverse datasets under a unified protocol.
Overall, Dataforge delivers the greatest and most consistent improvements across both classification and regression benchmarks. Compared with classical automated feature engineering and search methods (e.g., AutoFeat, AFAT, NFS, TTG, and OpenFE), Dataforge more reliably discovers beneficial transformations, suggesting that iterative, feedback-driven planning is better suited for navigating large operator spaces than one-shot heuristics or fixed search strategies. Compared with RL-based pipelines (GRFG/MOAT) and the Pure LLM variant, Dataforge exhibits substantially higher stability: it avoids low-utility or invalid proposals and maintains strong performance across heterogeneous datasets with different sizes and feature dimensions. These results indicate that coarse routing to an appropriate feature-operation family, followed by fine-grained action planning and grounding, effectively reduces candidate ambiguity and improves decision quality under a fixed budget.
Insights: First, the consistent improvements over Original features validate that Dataforge can autonomously refine tabular representations via a combination of selection, transformation, and feature construction, rather than relying on a single operator family. Second, the performance advantages are achieved without training a task-specific policy (unlike RL methods), making the approach readily deployable across datasets. Finally, Dataforge requires minimal user input (dataset and basic runtime constraints) while producing a fully automated pipeline, supporting the practicality of agentic data preparation as a general paradigm for tabular feature engineering.
4.3. Ablation Study
Table 2 ablates key components of Dataforge on three representative datasets (German Credit, Amazon Employee, and OpenML 618). For each variant, we report the downstream score (macro-F1 for classification and 1-RAE for regression), the valid-action rate Valid(%) (fraction of proposed actions that pass grounding and execute successfully), and the average number of evaluator calls, which reflects the interaction cost of the iterative pipeline.
Table 2. Ablation study of Dataforge on three representative datasets. Score is macro-F1 (%) for classification and 1-RAE for regression. Valid(%) reports the ratio of proposed actions that pass grounding and are executed successfully; Calls denotes the average number of evaluator calls.
German Credit (C) Amazon Employee (C) OpenML 618 (R)
Variant Score Valid(%) Calls Score Valid(%) Calls Score Valid(%) Calls
Full (Dataforge) 79.60% 98.51% 3 94.41% 99.90% 6 0.7219 98.84% 4
w/o Routing 76.86% 97.01% 2 93.96% 98.33% 3 0.7169 90.41% 3
w/o Grounding 77.49% 91.60% 5 94.36% 84.10% 11 0.7031 75.76% 5
w/o Perf Loop 76.27% 96.30% 1 93.95% 97.00% 1 0.6653 98.25% 1
[…]
Insights: These ablations highlight complementary roles of the components. Grounding primarily improves reliability by filtering non-executable actions, thereby reducing wasted attempts even when the final score is not dramatically affected. Hierarchical routing improves decision quality by narrowing the candidate space; without it, the agent is more likely to try low-utility actions and stop early under the same auto-stop rule. The performance optimization loop is crucial for refinement: iterative evaluation and best-so-far updates are needed to move beyond a first valid action and consistently reach the strongest pipeline under a fixed budget.
4.4. Study of Reliability and Automation.
We evaluate whether Dataforge can run end-to-end autonomously and remain reliable under iterative feature operations. In this study, we focus on two practical indicators: (i) execution reliability, measured by the rate of invalid attempts during action proposal and execution, and (ii) automation stability, reflected by the number of retries needed to obtain valid actions and complete a run.
(a) F1-score comparison.
(b) Error rate comparison.
Figure 3. Comparison of modular agent (DataAgent) and pure LLM baselines, showing higher reliability and fewer retries for the modular agent paradigm.
As shown in Figure 3, Dataforge exhibits consistently higher reliability than pure LLM baselines. Across all nine datasets, it completes the entire pipeline without interruption and typically converges with about two action retries on average. In contrast, directly using Llama 3.1 8B or Llama 3 8B to propose feature operations leads to a non-trivial failure rate (around 3–5%), indicating that purely generative pipelines are more likely to produce invalid or non-executable operations. Meanwhile, Dataforge achieves comparable or slightly better downstream utility (average score $\approx 0.783$), suggesting that improved reliability does not come at the expense of performance.
Insights: These results highlight the importance of the modular agent design. By combining hierarchical routing with grounding validation, Dataforge constrains the action space and filters out unsafe proposals before execution, thereby reducing retries and preventing workflow breakdowns. This makes the system more predictable for continuous operation and supports the goal of fully automated data preparation with minimal user intervention.
4.5. Study of Efficiency Trade-offs.
We further study the efficiency–performance trade-offs of Dataforge by comparing it with RL-based pipelines that explicitly optimize feature operations via training. We report three aspects of efficiency in Table 3: training cost (wall-clock training time), inference latency (time to complete a single run), and interaction cost measured by the number of calls (i.e., the number of evaluator/tool invocations needed to obtain a final result). This experiment aims to characterize whether an agentic pipeline can remain responsive while still achieving strong downstream utility.
Table 3. Efficiency comparison: time (seconds) and call times.
Metric RL-Policy-1 (GRFG (Wang et al., 2022)) RL-Policy-2 (MOAT (Wang et al., 2023)) Dataforge
Training Time (sec) 1493.7 649.4 0
Inference Time (sec) 0.5 2.14 3.89
Calls 13 22 2
Performance 0.5768 0.6251 0.7832
Table 3 shows that Dataforge eliminates the dominant overhead of RL-style optimization. Both RL-Policy-1 (GRFG) and RL-Policy-2 (MOAT) require substantial training time (1493.7 s and 649.4 s), whereas Dataforge shows zero training cost and can be directly applied to new datasets. At inference time, Dataforge completes a full workflow in 3.89 s on average, which is slightly slower than RL inference (0.5–2.14 s) but remains lightweight in absolute terms. Notably, it achieves this with only two calls on average, compared to 13–22 calls for the RL baselines, indicating a markedly lower interaction/sample complexity under the same evaluation setting. Despite the reduced call budget, Dataforge attains the best downstream performance (0.7832), suggesting that its iterative planning is highly cost-effective.
Insights: These results highlight a favorable operating point for practical deployment. While RL policies can achieve fast inference after training, the up-front training time and higher call counts make them less attractive when datasets change frequently or when rapid, on-demand data preparation is required. In contrast, Dataforge trades a small increase in per-run latency for eliminating training altogether and substantially reducing interaction cost, yielding strong performance with high responsiveness. This supports the use of modular agentic feature engineering as a practical alternative to training-heavy optimization methods.
5. Demo and Case Studies
5.1. User Interface
Figure 4. The interface of Dataforge.
Figure. 4 illustrates the user interface of Dataforge, which offers an intuitive environment for autonomous tabular data transformation. Users can upload datasets in standard formats (.csv, .parquet, and .hdf), label predictors and response column in a machine learning task, and configure runtime parameters, such as, Max Actions, Min Improvement, and Time Budget. After uploading datasets, users can select from pre-loaded examples. During execution, the interface visualizes real-time logs of thinking steps in reasoning, data cleaning and feature engineering actions, and predictive performance metrics (e.g., F1-score, AUC, MAE, and RMSE depending on classification, regression, and unsupervised tasks ), so that users can observe how the data agent progressively refines data quality. Upon completion, it summarizes all executed actions, validation outcomes, and final performance statistics, to provide transparency and provenance, as well as collect action historical trajectories for benchmark data of training data agents. The interface is designed to be modular and extensible, supporting future integration with interactive dashboards and collaborative agent environments.
5.2. Toward Accurate and Efficient Heart-Disease Detection
@@
Figure 5. Dataforge Dealing with Heart-Disease Detection.
@@
Throughout this process (Figure. 5), no manual intervention or algorithmic knowledge is needed. The workflow completes in about 100 seconds and generates a concise report and a new dataset that achieves higher diagnostic accuracy with fewer features, reducing the total number of variables from 44 to 20 while improving the predictive score from 0.772 to 0.840. This demonstrates that DataForge can serve as an intelligent assistant for cardiology, automatically transforming complex measurements into a cleaner, smaller, and more reliable dataset for accurate and efficient heart-disease detection.
5.3. Chemical Design
Figure 6. Dataforge for materials discovery.
Figure 6 illustrates how Dataforge supports a practical materials discovery workflow. Starting from a small seed set of physical experiments, we obtain an initial tabular dataset with multiple target properties. Dataforge then enhances the dataset via automated feature operations and produces an AI-ready version together with traceable logs and a natural-language report. Training a downstream property predictor on the enhanced data improves predictive utility across targets, which enables more reliable screening and prioritization of candidate formulations, thereby reducing the number of costly follow-up experiments.
6. Related Work
6.1. Automated Feature Engineering
Feature transformation. Prior work formulates feature transformation as searching over composition trajectories of operators and operands. Early systems suggest generating a large candidate pool followed by selection to control redundancy (Katz et al., 2016). To improve trajectory quality, many methods adopt structured discrete search, including heuristic-guided exploration and other discrete optimization strategies (Kanter and Veeramachaneni, 2015; Khurana et al., 2016; Tran et al., 2016), as well as evolutionary algorithms for population-based exploration (Zhu et al., 2022a; Gong et al., 2024). Reinforcement learning further models transformations as sequential decision making and learns policies from downstream rewards (Wang et al., 2022; Ying et al., 2025b; Bai et al., 2025b), with traceable/evolutionary variants emphasizing controllable search dynamics and interpretability (Xiao et al., 2023a, 2024). Beyond discrete search, continuous optimization methods embed feature programs into differentiable latent spaces and apply gradient-based updates (Wang et al., 2023; Zhu et al., 2022b; Gong et al., 2025b), while facing the challenge of mapping optimized embeddings back to valid executable transformations.
Feature selection. Classical feature selection includes similarity/spectral criteria (e.g., Laplacian Score (He et al., 2005), SPEC (Zhao and Liu, 2007), Fisher Score (Duda et al., 2006), Trace Ratio (Nie et al., 2010)), information-theoretic filters (e.g., Information Gain (Lewis, 1992), mutual information (Battiti, 1994), mRMR (Peng et al., 2005), FCBF (Yu and Liu, 2003)), sparse learning (e.g., $\ell_{p}$-regularized methods (Tibshirani, 1996; Zhu et al., 2003)), and statistics-based measures (e.g., T-test (Davis et al., 1986), Chi-square (Liu and Setiono, 1995), Gini (Gini, 1912), CFS (Hall and Smith, 1999)). Recent work leverages deep and RL formulations to improve scalability and search quality, including deep feature selection (Xiao et al., 2023b; Ying et al., 2024a, b), RL-based selection (Wang et al., 2025e), Monte-Carlo style methods (MARFS (Liu et al., 2023), MCRFS (Liu et al., 2021b)), and specialized settings such as incomplete multi-view unsupervised selection (Huang et al., 2023).
LLMs for feature engineering. Recent studies explore using LLMs to propose feature transformations or engineered features via natural-language reasoning, enabling broader operator coverage and reduced manual effort (Han et al., 2024; Hollmann et al., 2023; Gong et al., 2025c, d, a; Wang et al., 2025c; Xie et al., 2025). However, purely generative pipelines can suggest invalid or non-executable actions and often require additional validation and iterative control to ensure reliability and stable improvements.
6.2. Agentic Platform
Agentic systems augment LLMs with multi-step reasoning, planning, and tool use (Zheng et al., 2026; Wang et al., 2025a; Xie and Ju, 2026), enabling them to iteratively act in an environment rather than producing one-shot outputs (Yin et al., 2025). A representative paradigm is ReAct, which interleaves reasoning traces with action/tool calls to improve task completion and error recovery (Yao et al., 2022).
Building on general tool-using agents, web agents study how LLMs perceive and interact with complex web environments (e.g., multi-page, multi-tab interfaces) to accomplish real-world browsing tasks. Benchmarks such as WebArena provide realistic web interaction environments for evaluating such agents (Zhou et al., 2023), and systems like WebVoyager demonstrate end-to-end web navigation and action execution at scale (He et al., 2024).
Recently, the same agentic principles have been extended to the data domain. Fu et al. propose autonomous data agents that combine task decomposition, action reasoning, grounding into executable code/tool calls, and adaptive workflow planning to automate data-to-knowledge pipelines (Fu et al., 2025). This perspective highlights new opportunities and challenges for trustworthy, scalable data operations (e.g., preprocessing, selection, transformation, and repairs) and motivates agentic platforms like ours that emphasize grounded execution and iterative feedback for reliable data engineering (Fu et al., 2025).
7. Conclusion
We presented Dataforge, an LLM-powered agentic data engineering platform for tabular data. It is automatic, safe, and non-expert friendly. By combining hierarchical routing with evaluation-driven refinement and automatic stopping, Dataforge transforms raw datasets into AI-ready representations end-to-end. Experiments on nine public benchmarks show strong and robust gains over classical AutoFE/search, RL-based, and LLM baselines, while ablations highlight the roles of routing, grounding, and iterative optimization.